• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 00:01:18 on March 6,2006

    Entry number 165412

    keyword=Shift Summary

    Sunday March 5, 2006 swing shift
    J. LeRose (SL)
    G. M. Urciuoli (TO)
    F. Garibaldi

    Program is take production data at 5 micro amps
    Beam energy 1519 MeV

    Note from previous shift:
    observing: Time diff of helicity to nonhelicity times = 501.96949 seconds
    In end of run is worrisome. They tried rebooting ROC 23 (see halog 165354)
    17:10 beam off for target polarization measurement (39.6% Halog165364)
    helicity time difference is still large (625) in run 2677. Bob Michaels is looking into it. See Halog 165365. We will reboot ROC23 again after the next run.
    17:57 rebooted roc 23 also reset prescales ps1 3300, ps2 90
    18:00 BB chambers tripped again
    18:40 Beam 1/2 wave goes IN
    21:08 beam off for target polarization measurement (40.7% halog 165398)
    21:16 beam back
    22:08 Bob Michaels calls in to say that he thinks the helicity difference issue is a false alarm. He sees no problem with the actual data. There's only a problem with the online vme code. See halog 165401. He suggests that others check his conclusion and says if time permits he will try to fix/debug the online vme code tomorrow.
    23:55 ran QASY on run 2684 got essentially the same results as in the end of run log entry (Beam charge asymmetry ~100 ppm)

    Run start/stop events/comments
    2677 16:01/17:09 1M helicity time diff=625 (still not good)
    2678 17:24/17:53 0.332 M stopped for HV trip BB chambers (S13)
    rebooted ROC 23
    2679 junk
    2680 17:57 more junk need to reset prescales
    2681 18:02/18:39 .357 M stopped for change of beam 1/2 wave plate from Out to IN. Helicity time difference is still large.
    2682 18:41/19:54 1M Helicity time difference is good now!
    2683 19:55/21:08 1M
    2684 21:16/22:39 BB HV trip during run, N HV too. Stopped when we noticed the n-arm had tripped.
    2685 22:39/23:46 1M
    2686 22:50/ in progress at end of shift